# Bridging Quantum Mechanics to Organic Liquid Properties via a Universal Force Field

### v3

@@ 1.4 Applications in Liquid Electrolytes
One important application of organic liquid simulation is the design of electrolytes for batteries. In addition to thermodynamic properties, transport properties such as viscosity, conductivity, and self-diffusion coefficient are also critical for electrolytes. For typical electrolyte systems, ByteFF-Pol is compared with two conventional force fields, OPLS-AA and OPLS-AA-SC. Instead of the original OPLS-AA charges, OPLS-AA-SC employs CHELPG charges [6] scaled by a factor of 0.8, an empirical charge model widely used to approximate polarization in electrolyte systems.
Figure 4: Evaluation of ByteFF-Pol for predicting electrolyte properties. (a-c) Comparison of different force fields in predicting (a) density, (b) viscosity and (c) conductivity of liquid electrolytes, respectively. Pearson correlation coefficients (r) of each force field are provided in the legends. (d-e) Examples for conductivity at (d) varying lithium salt concentrations; (e) at different temperatures (solid lines denote simulation results, dashed lines denote experimental data). The numbers in the legends indicate the approximate molar fraction of each species. (f) Examples for self-diffusion coefficient ($D$) of different species in electrolyte. Upper bars correspond to $D$ of species X (labeled on the x axis) in the X:EMC binary solvent (molar ratio=1:1); and lower bars show the $D$ of X in the LiPF6:X:EMC electrolytes (molar ratio=1:4:4). Experimental data are obtained from [45]. The full names and SMILES strings corresponding to the molecular abbreviations are provided in Supplementary Table 1.
@@
OPLS-AA 88 0.022 1.76% 0.98 91 12.20 163.3% 0.77
OPLS-AA-SC 88 0.054 4.38% 0.97 91 1.79 39.8% 0.83
ByteFF-Pol 88 0.024 1.94% 0.99 91 1.06 20.0% 0.96
@@
As shown in Figure 4 (a, b, and c), ByteFF-Pol exhibits excellent accuracy in predicting the density, viscosity, and conductivity of electrolyte systems, outperforming other evaluated force fields. OPLS-AA maintains high accuracy for density but overestimates viscosity and underestimates conductivity. While the scaled charge approach employed by OPLS-AA-SC improves conductivity accuracy, it introduces underestimation of both density and viscosity. In contrast, ByteFF-Pol accurately predicts all three properties, emphasizing the importance of explicitly modeling polarization. On this extensive conductivity benchmark encompassing nearly 5,000 electrolyte systems with diverse components and temperature conditions, ByteFF-Pol achieves a Pearson correlation coefficient of 0.95, demonstrating its utility as a tool for screening high-conductivity electrolytes. For other electrolyte force fields, large-scale benchmarking of this scope was not feasible; however, comparative analyses provided in Supplementary Section 9 underscore ByteFF-Pol’s advantages.
To further demonstrate the utility of ByteFF-Pol in electrolyte design, we present representative examples highlighting its practical potential. In real-world research, a key objective is to identify the optimal lithium salt concentration that maximizes electrolyte conductivity—a goal achievable by leveraging ByteFF-Pol to predict conductivity across varying lithium salt concentrations. Examples encompass systems containing two types of lithium salts, paired with cyclic carbonates, linear carbonates, esters, ethers, and nitriles, with conductivity values spanning 0.1 to $35\text{\,}\mathrm{mS}\text{\,}{\mathrm{cm}}^{-1}$ are provided in Figure 4 (d). It is shown that ByteFF-Pol not only delivers accurate conductivity predictions for diverse systems but also precisely captures the positions of conductivity peaks.
@@
The diffusion coefficient of solvent molecules in electrolytes is a key property linked to the solvation structure of lithium ions [45]. Given that the definition and derivation of solvation structures lack clear standardization, we opted to directly benchmark ByteFF-Pol’s performance by comparing its predicted diffusion coefficients against experimental measurements. Figure 4(f) presents these coefficients for various solvents in systems with and without lithium salts. For common esters and carbonates, ByteFF-Pol’s predictions exhibit excellent agreement with experimental results. For fluorinated solvents, ByteFF-Pol shows a systematic underestimation, which, as discussed previously, can be attributed to its limitations in capturing anisotropic interactions. Despite this, ByteFF-Pol correctly predicts the experimental trends even for these challenging systems. In comparison, OPLS-AA consistently underestimates the diffusion coefficient across all systems, while OPLS-AA-SC consistently overestimates it.
In summary, ByteFF-Pol exhibits excellent accuracy in predicting diverse electrolyte properties. This capability, on one hand, validates the feasibility of developing universal force fields from ab initio calculations and, on the other hand, paves the way for exploring novel electrolytes across a vast chemical space.
@@ 3.1 ALMO-EDA
This interaction energy is partitioned into physically meaningful components using the second-generation ALMO-EDA [28]. This variational scheme decomposes the total interaction energy as follows:
@@
The frozen term is further decomposed into dispersion ($U_{\mathrm{disp}}^{\mathrm{EDA}}$), electrostatic ($U_{\mathrm{est}}^{\mathrm{EDA}}$), and Pauli ($U_{\mathrm{pauli}}^{\mathrm{EDA}}$) terms [21]:
@@ 3.2 Force Field Formulation
The non-bonded energy, $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$, comprises five terms:
@@
Repulsion and dispersion interactions are described by a modified Buckingham (exp-6) potential [38, 5], which has shown superior agreement with DFT-computed intermolecular energies compared to the Lennard-Jones potential [42, 5]. The functional forms are:
@@
Here, $r_{ij}$ is the interatomic distance between atoms $i$ and $j$, $\epsilon^{\mathrm{rep}}$, $\lambda^{\mathrm{rep}}$, $C_{6}$, and $r^{*}$ are atom-specific parameters predicted by the GNN model. An additional $1/r^{12}$ term in the repulsion potential, scaled by a small constant $s_{\mathrm{rep}}$ (1.5 $\cdot(\mathrm{kcal/mol})^{1/12}$), is included to prevent unphysical attraction at very short interatomic distances. The dispersion term is damped at short ranges to avoid singularities and the double-counting of correlation effects, analogous to the DFT-D3 treatment [18]. The dimensionless damping factor $s_{\mathrm{disp}}$ is set to 120. As demonstrated in Section 1.2, this damped dispersion function provides an excellent fit to the target ALMO-EDA data.
The electrostatic interaction are modeled following the AMOEBA polarizable framework [11]. We use atom-centered point charges for the permanent electrostatics and an induced dipole model for polarization:
@@
This simplification from a full multipole expansion to monopoles streamlines parameterization without a significant loss of accuracy. The induced dipole, $\boldsymbol{\mu}^{\mathrm{ind}}$, is determined self-consistently by solving the polarization equations in response to the electric field $\mathbf{E}$. For efficient training, $\boldsymbol{\mu}^{\mathrm{ind}}$ is computed via direct matrix inversion, as detailed in Supplementary Section 6. The atomic charge $q$ and polarizability alphaalpha used in solving $\mu^{\mathrm{ind}}$ are predicted by the GNN model.
Finally, the charge-transfer term is modeled using a functional form inspired by the work of Deng et al. [12]:
@@
The atomic non-bonded parameters are predicted by the GNN model through several distinct approaches. Similar to our previous work on ByteFF, atomic charges $q$ are predicted using a Bond Charge Correction (BCC) model to preserve charge conservation. Furthermore, a subset of parameters are obtained using a physically-inspired Tkatchenko-Scheffler (TS) scaling model [48]:
@@
Once all atom-wise parameters are determined, the pairwise interaction parameters are derived using combination rules. We employ the standard Lorentz-Berthelot (LB) rules [36], which use an arithmetic mean for $r_{i}^{*}$ and a geometric mean for all other parameters:
@@ 3.3 Dataset
ByteFF-Pol’s training relies on three core datasets: the existing optimization and torsion datasets from our previous work and a newly created dimer dataset. Details on the original optimization and torsion datasets are in our previous publication [55]. For this work, we’ve augmented the optimization dataset by calculating the minimum basis iterative stockholder (MBIS) charges and volumes for each molecule at the PBE0/def2-TZVPD level, which are used to pre-train the model.
@@
To benchmark the performance of ByteFF-Pol, we used five datasets of experimental physical properties. The first contains 96 pure molecular liquids with known densities and evaporation enthalpies from the QUBEKit benchmark [23]. The second and third datasets consist of experimental densities for 595 binary liquids from the NIST ThermoML archive [8] and 2132 pure molecular liquids from the CRC Handbook [35], respectively. For electrolyte-relevant systems, we compiled two additional datasets collected from literatures [54]: one with 88 experimental densities and 91 experimental viscosities for pure solvents and lithium electrolytes, and another containing 4896 conductivity measurements for various lithium electrolytes across different compositions and temperatures. The number of results for each force field may differ from that of the benchmark set, owing to parametrization errors or MD simulation failures.
@@ 3.4 Training Techniques
In the initial pre-training stage, the GNN model is trained to reproduce the bonded parameters from ByteFF [55], along with atomic charges and volumes from MBIS calculations. This stage allows the model to develop atomic hidden states that recognize the chemical environment and provides a physically-grounded starting point for the subsequent training.
In the main training stage, the 6 million data points are split into 5.4 million for training and 0.6 million for validation. We freeze most of the model and train only the output layers corresponding to the non-bonded interaction parameters. The training targets are the total interaction energies and their decomposed components from the ALMO-EDA calculations. Our force field’s energy decomposition scheme is carefully designed to align with the ALMO-EDA scheme, ensuring that the model learns physically meaningful interactions, as detailed in Supplementary Section 7.
Finally, in the fine-tuning stage, we update the torsion parameters to ensure they are consistent with the newly trained non-bonded interactions. The training target for this stage is the set of torsion scan energy profiles from ByteFF torsion dataset [55].
@@ 4 Data availability
The source data for Figure 3 and Figure 4 are provided in the supplementary information. A subset of the training dataset used in this study will be released with the source code.
@@ 5 Code availability
The code for ByteFF-Pol including the training scheme and trained model will be released upon acceptance of this paper.

### v4

@@ 1.4 Applications in Liquid Electrolytes
One important application of organic liquid simulation is the design of electrolytes for batteries. In addition to thermodynamic properties, transport properties such as viscosity, conductivity, and self-diffusion coefficient are also critical for electrolytes. For typical electrolyte systems, ByteFF-Pol is compared with two conventional force fields, OPLS-AA and OPLS-AA-SC, as well as a SOTA polarizable force field, AMOEBA. Instead of the original OPLS-AA charges, OPLS-AA-SC employs CHELPG charges [breneman1990determining] scaled by a factor of 0.8, an empirical charge model widely used to approximate polarization in electrolyte systems.
Figure 4: Evaluation of ByteFF-Pol for predicting electrolyte properties. (a-c) Comparison of different force fields in predicting (a) density, (b) viscosity and (c) conductivity of liquid electrolytes, respectively. Pearson correlation coefficients (r) of each force field are provided in the legends. (d-e) Examples for conductivity at (d) varying lithium salt concentrations; (e) at different temperatures (solid lines denote simulation results, dashed lines denote experimental data). The numbers in the legends indicate the approximate molar fraction of each species. (f) Examples for self-diffusion coefficient ($D$) of different species in electrolyte. Upper bars correspond to $D$ of species X (labeled on the x axis) in the X:EMC binary solvent (molar ratio=1:1); and lower bars show the $D$ of X in the LiPF6:X:EMC electrolytes (molar ratio=1:4:4). Experimental data are obtained from [C9EE00141G]. The full names and SMILES strings corresponding to the molecular abbreviations are provided in Supplementary Table 1.
@@
OPLS-AA 88 0.022 1.8% 0.98 91 12.20 163.3% 0.77
OPLS-AA-SC 88 0.054 4.4% 0.97 91 1.79 39.8% 0.83
AMOEBA 88 0.030 2.4% 0.98 90 6.69 103.4% 0.79
ByteFF-Pol 88 0.024 1.9% 0.99 91 1.06 20.0% 0.96
@@
As shown in Figure 4, ByteFF-Pol exhibits excellent accuracy in predicting the density, viscosity, and conductivity of electrolyte systems, outperforming other evaluated force fields. OPLS-AA maintains high accuracy for density but overestimates viscosity and underestimates conductivity. While the scaled charge approach employed by OPLS-AA-SC improves conductivity accuracy, it introduces underestimation of both density and viscosity. These discrepancies primarily arise from the neglect of explicit polarization in conventional force fields, a factor demonstrated to be critical for electrolyte systems in the ablation study (Supplementary Section 11). Although AMOEBA achieves better alignment with experimental reference by accounting for polarization, it still exhibits a systematic overestimation of viscosity and underestimation of conductivity. This limitation might be attributed to its training protocol, which relies on fitting to experimental bulk data; consequently, the force field’s performance may diminish when the chemical space of the electrolyte systems extends beyond the coverage of its original experimental training set. Furthermore, AMOEBA suffers from occasional parameterization failures and simulation instabilities, resulting in unavailable conductivity data for approximately one-quarter of the tested systems. In contrast, ByteFF-Pol accurately predicts all three properties, emphasizing the importance of explicitly modeling polarization. On this extensive conductivity benchmark encompassing nearly 5,000 electrolyte systems with diverse components and temperature conditions, ByteFF-Pol achieves a Pearson correlation coefficient of 0.95, while yielding the lowest MAE and MAPE, demonstrating its utility as a tool for screening high-conductivity electrolytes. For other electrolyte force fields, large-scale benchmarking of this scope was not feasible; however, comparative analyses provided in Supplementary Section 9.2 underscore ByteFF-Pol’s advantages. As shown in Supplementary Figure 15 and Supplementary Table 11, ByteFF-Pol exhibits superior accuracy in predicting conductivity compared to BAMBOO, the current SOTA machine learning force field for electrolyte systems. The experimental conductivity data presented in Supplementary Figure 15 were obtained from iterative optimization of an electrolyte system [zhu2024differentiable]. As a result of the excellent agreement between ByteFF-Pol’s predictions and experimental observations, the model holds significant potential to replace such experimental measurements, thereby reducing associated costs.
To further demonstrate the utility of ByteFF-Pol in electrolyte design, we present representative examples highlighting its practical potential. In real-world research, a key objective is to identify the optimal lithium salt concentration that maximizes electrolyte conductivity—a goal achievable by leveraging ByteFF-Pol to predict conductivity across varying lithium salt concentrations. Examples encompass systems containing two types of lithium salts, paired with cyclic carbonates, linear carbonates, esters, ethers, and nitriles, with conductivity values spanning 0.1 to 35 mS/cm are provided in Figure 4 (d). It is shown that ByteFF-Pol not only delivers accurate conductivity predictions for diverse systems but also precisely captures the positions of conductivity peaks.
@@
The diffusion coefficient of solvent molecules in electrolytes is a key property linked to the solvation structure of lithium ions [C9EE00141G]. Given that the definition and derivation of solvation structures lack clear standardization, we opted to directly benchmark ByteFF-Pol’s performance by comparing its predicted diffusion coefficients against experimental measurements. Figure 4(f) presents these coefficients for various solvents in systems with and without lithium salts. For common esters and carbonates, ByteFF-Pol’s predictions exhibit excellent agreement with experimental results. For fluorinated solvents, ByteFF-Pol shows a systematic underestimation, which, as discussed previously, can be attributed to its limitations in capturing anisotropic interactions. Despite this, ByteFF-Pol correctly predicts the experimental trends even for these challenging systems. In comparison, OPLS-AA consistently underestimates the diffusion coefficient across all systems, while OPLS-AA-SC consistently overestimates it.
Owing to its relatively simple functional form, ByteFF-Pol achieves a simulation speed of approximately 40 ns/day for a 10,000-atom system on an A100 GPU, and 32 ns/day on the more cost-effective A10 GPU. This computational efficiency is markedly superior to that of typical ML force fields (see Supplementary Section 9.3 for detailed comparison). In summary, ByteFF-Pol exhibits excellent accuracy and efficiency in predicting diverse electrolyte properties. This capability, on one hand, validates the feasibility of developing universal force fields from ab initio calculations and, on the other hand, paves the way for exploring novel electrolytes across a vast chemical space.
@@ 3.1 ALMO-EDA
This interaction energy is partitioned into physically meaningful components using the second-generation ALMO-EDA [khaliulinEDAv1Original2007]. This variational scheme decomposes the total interaction energy as follows:
@@
The frozen term is further decomposed into dispersion ($U_{\mathrm{disp}}^{\mathrm{EDA}}$), electrostatic ($U_{\mathrm{est}}^{\mathrm{EDA}}$), and Pauli ($U_{\mathrm{pauli}}^{\mathrm{EDA}}$) terms [hornEDAElectrostaticDispersion2016]:
@@ 3.2 Force Field Formulation
The detailed bonded terms could be found in Supplementary Section 6.2. The non-bonded energy, $U_{\mathrm{non-bonded}}^{\mathrm{FF}}$, comprises five terms:
@@
Repulsion and dispersion interactions are described by a modified Buckingham (exp-6) potential [mason_transport_1954, borodin_development_2006], which has shown superior agreement with DFT-computed intermolecular energies compared to the Lennard-Jones potential [rowley_determination_2001, borodin_development_2006]. The functional forms are:
@@
Here, $r_{ij}$ is the interatomic distance between atoms $i$ and $j$, $\epsilon^{\mathrm{rep}}$, $\lambda^{\mathrm{rep}}$, $C_{6}$, and $r^{*}$ are atom-specific parameters predicted by the GNN model. An additional $1/r^{12}$ term in the repulsion potential, scaled by a small constant $s_{\mathrm{rep}}$ (1.5 $\cdot(\mathrm{kcal/mol})^{1/12}$), is included to prevent unphysical attraction at very short interatomic distances. The dispersion term is damped at short ranges to avoid singularities and the double-counting of correlation effects, analogous to the DFT-D3 treatment [grimme_consistent_2010]. The dimensionless damping factor $s_{\mathrm{disp}}$ is set to 120. As demonstrated in Section 1.2, this damped dispersion function provides an excellent fit to the target ALMO-EDA data.
The electrostatic interaction are modeled following the AMOEBA polarizable framework [demerdashAssessingManybodyContributions2017]. We use atom-centered point charges for the permanent electrostatics and an induced dipole model for polarization:
@@
This simplification from a full multipole expansion to monopoles streamlines parameterization without a significant loss of accuracy. The induced dipole, $\boldsymbol{\mu}^{\mathrm{ind}}$, is determined self-consistently by solving the polarization equations in response to the electric field $\mathbf{E}$. For efficient training, $\boldsymbol{\mu}^{\mathrm{ind}}$ is computed via direct matrix inversion, as detailed in Supplementary Section 6. The atomic charge $q$ and polarizability $\alpha$ used in solving $\mu^{\mathrm{ind}}$ are predicted by the GNN model.
Finally, the charge-transfer term is modeled using a functional form inspired by the work of Deng et al. [deng2017estimating]:
@@
The atomic non-bonded parameters are predicted by the GNN model through several distinct approaches. Similar to our previous work on ByteFF, atomic charges $q$ are predicted using a Bond Charge Correction (BCC) model to preserve charge conservation. Furthermore, a subset of parameters are obtained using a physically-inspired Tkatchenko-Scheffler (TS) scaling model [tkatchenko_accurate_2009]:
@@
Once all atom-wise parameters are determined, the pairwise interaction parameters are derived using combination rules. We employ the standard Lorentz-Berthelot (LB) rules [lorentz_ueber_1881], which use an arithmetic mean for $r_{i}^{*}$ and a geometric mean for all other parameters:
@@ 3.3 Dataset
ByteFF-Pol’s training relies on three core datasets: the existing optimization and torsion datasets from our previous work and a newly created dimer dataset. Details on the original optimization and torsion datasets are in our previous publication [zheng2025data]. For this work, we’ve augmented the optimization dataset by calculating the minimal basis iterative Stockholder (MBIS) charges and volumes for each molecule at the PBE0/def2-TZVPD level, which are used to pre-train the model.
@@
To benchmark the performance of ByteFF-Pol, we used five datasets of experimental physical properties. The first contains 96 pure molecular liquids with known densities and evaporation enthalpies from the QUBEKit benchmark [hortonQUBEKitAutomatingDerivation2019]. The second and third datasets consist of experimental densities for 595 binary liquids from the NIST ThermoML archive [chirico2003thermoml] and 2132 pure molecular liquids from the CRC Handbook [Lide2004CRC], respectively. For electrolyte-relevant systems, we compiled two additional datasets collected from literature [yang2025unified]: one with 88 experimental densities and 91 experimental viscosities for pure solvents and lithium electrolytes, and another containing 4896 conductivity measurements for various lithium electrolytes across different compositions and temperatures. The number of results for each force field may differ from that of the benchmark set, owing to parametrization errors or MD simulation failures.
@@ 3.4 Training Techniques
In the initial pre-training stage, the GNN model is trained to reproduce the bonded parameters from ByteFF [zheng2025data], along with atomic charges and volumes from MBIS calculations. This stage allows the model to develop atomic hidden states that recognize the chemical environment and provides a physically-grounded starting point for the subsequent training.
In the main training stage, the 6 million data points are split into 5.4 million for training and 0.6 million for validation. We freeze most of the model and train only the output layers corresponding to the non-bonded interaction parameters. The training targets are the total interaction energies and their decomposed components from the ALMO-EDA calculations. We employ a scaled mean squared error (SMSE) loss for train each energy term:
$\mathcal{L}_{\mathrm{SMSE}}^{\mathrm{term}}=\frac{1}{N}\sum_{i}s_{i}^{\mathrm{Boltz}}s_{i}^{\mathrm{Force}}\left(U^{\mathrm{FF,inter}}_{\mathrm{term},i}-U^{\mathrm{EDA}}_{\mathrm{term},i}\right)^{2},$ (16)
where “term” refers to a specific interaction term, i.e., “dispersion”, “electrostatic with Pauli repulsion”, “polarization” and “charge transfer”. The two scaling factors, $s_{i}^{\mathrm{Boltz}}$ and $s_{i}^{\mathrm{Force}}$, reduce the weighting of data points with large interaction energies and large interaction forces, respectively. This ensures a stable training process and improves the accuracy for conformations that have higher Boltzmann sampling probabilities. Detailed definitions of the scaling factors and the decomposed energy terms are provided in Supplementary Section 7.2.
Finally, in the fine-tuning stage, we update the torsion parameters to ensure they are consistent with the newly trained non-bonded interactions. The training target for this stage is the set of torsion scan energy profiles from ByteFF torsion dataset [zheng2025data].
@@ 4 Data availability
A subset of the training dataset and a subset of the validation dataset that is used to reproduce the results in Figure 2 (a) are released with the source code. Source Data for Figure 2, 3 and 4 are provided with this paper.
@@ 5 Code availability
The code for ByteFF-Pol including the training scheme, trained model and examples are available at https://github.com/ByteDance-Seed/byteff2.